myininaya:

Find the principal square root of 251001 without a calculator.

moreso:

501

myininaya:

Did you get that without a calculator?

myininaya:

Oh okay... I made up the question purposely so if would be factorable like a quadratic in terms of the base 10^2. I thought it would be more fun but I guess not :p \[ 251001 \\ 25 \cdot 10^4+10 \cdot 10^2 +1 \\ (5 \cdot 10^2+1)^2 \\ (500+1)^2 \\ (501)^2 \]

moreso:

Oh cool. I did not see that.

myininaya:

Yea I like it... :) Like factor 156 in a similar way 1*10^2+5*10+6 (10+3)(10+2) (13)(12) yes I know this doesn't work for all numbers but it is fun for the numbers it works for

myininaya:

I wrote 156 as a quadratic expression in terms of 10

myininaya:

Numbers in the form abc where a, b, and c are digits of the number abc can be factored like this if we can find nonnegative integers that multiplied to be a*c and also add up to be b So we know numbers like 156 143 176 275 and so on can be factored in this way

myininaya:

so factoring 300090006 should be doable without a calculator...

moreso:

what are the factors of that

myininaya:

\[300090006 \\ 3 00000000+90000+6 \\ 3 \cdot 10^8+9 \cdot 10^4+6 \\ 3 \cdot 10^8+3 \cdot 10^4+6 \cdot 10^4+6 \\ \text{ factor by grouping } \\ 3 \cdot 10^4(10^4+1)+6(10^4+1) \\ (3 \cdot 10^4+6)(10^4+1)\]

myininaya:

\[(30006)(10001)\]

myininaya:

that can be factored more

myininaya:

but it has been factored :p just not completely

myininaya:

\[3(10002)(10001) \\ 3(3)(334)(10001)\]

moreso:

Okay. And by factoring this way we can find if it's 'square root-able' , quickly.

myininaya:

oops not enough 3's in 334 3(3)(3334)(10001)

moreso:

clearly this number is not a perfect square, since $$300090006 = (3 \cdot 10^4+6)(10^4+1)$$

moreso:

Nevermind, that is wrong :)

myininaya:

hmmm I womder if we can say that... like let n be a positive integer \[(a \cdot 10^n+b)(c \cdot 10^n+e)\] I wonder if this is ever a perfect square besides when having the following case: a=c and b=e

myininaya:

I honestly don't know enough about number theory to know this

moreso:

A prime number will have no factoring.

myininaya:

I know ganeishie and some others and me talked about this a long time ago on openstudy but i can't remember if we talked about that

moreso:

Quickly factor this number: 100003

moreso:

using your method

myininaya:

\[100003 \\ 10 \cdot 10^4+3 \\ \text{ hmmm } .... 1 \cdot 10^5+3 \] not looking good

myininaya:

but maybe i should think about thinking in terms of something other than 10

moreso:

are you familiar with the 'brute force' way to factor a number. Algorithm: start with the smallest prime number 2, divide number by it, then keep dividing 2 until it is no longer divisible by 2. Next prime is 3, keep dividing number by 3...

moreso:

251001 clearly not divisible by 2 next we try 3.

myininaya:

ok back

myininaya:

the answer is yes

moreso:

251001 clearly not divisible by 2 next we try 3. oh it divides by 3 251001 /3 = 83667 divide the quotient by 3 83667 / 3 = 27889

myininaya:

did you know @moreso that you can tell if a number is divisible by 3 by adding up the digits of that number and seeing if that sum is divisible by 3 *if yes then the original number is divisible by 3 * if no then the original number is not divisible by 3

moreso:

251001 does it divide by 2 clearly not divisible by 2 since it has an odd number in the units place. Next we try 3. Is it divisible by 3? yes it divides by 3 251001 /3 = 83667 can we divide the quotient by 3 ? yes we can. 83667 / 3 = 27889 27889 /3 = 9296.3333 , no good. now we move on to 5, but the number ends in a 9. so don't check. move on to 7

moreso:

Right . There are many tricks.

myininaya:

yes and tricks are fun to explore

moreso:

With a calculator a number is divisible when the decimal place is 0, a crude test.

myininaya:

were you an OS user?

moreso:

This method is like a sieve. it is picking up all the factors.

moreso:

I used os a few times. Yes.

myininaya:

ah.. what was your username?

moreso:

benjc9

moreso:

I stopped using it a year ago.

myininaya:

well dang yep i don't remember you :(

myininaya:

have you studied any algorithms in number theory i think there are a few there

myininaya:

i think i remember something about pollard's something

moreso:

I was sneaky, i gave you a prime number to factor.

myininaya:

lol yep i used wolfram after my quadratic trick was not working

myininaya:

i can't remember it well

myininaya:

i remember they used to give lots of money for factoring really really really crazy large numbers

myininaya:

they did this i think to keep improving the security of like technology stuff

moreso:

Oh but why does factoring make it more secure? wouldnt it make it unsecure

myininaya:

if you are able to factor it then it is not secure so when someone is able to factor i think they up the amount of digits to make it more secure or something like that

moreso:

I see, but they aren't factoring by hand. They are using computers. Essentially these math people are programming algorithms.

myininaya:

read those first 2 sentences there

myininaya:

yep but they are creating factoring algorithms that computers can use to factor the numbers

myininaya:

oh didn't read your whole sentence

moreso:

Right.

myininaya:

now i can probably type up some algorithm for my factoring method above but my factoring method has a big chance of failing

myininaya:

for a lot of numbers

myininaya:

i would definitely not win lol

moreso:

if you can show how your quadratic method covers all possible cases of factoring then you have created , indirectly, a new primality test

moreso:

Because if a number cannot be factored quadratically, it has no factorization. Thus it is prime.

myininaya:

i believe not for sure that there are some numbers not in a quadratic form but perhaps some other polynomial form that can be factored maybe

moreso:

Also it seems to work better for larger number. For small numbers it is hard to factor quadratically 24 = 4 * 6 = 2^2 * 6

myininaya:

like 11011

myininaya:

i haven't really looked for a quadratic form here one sec

myininaya:

but i was thinking of 11*100+11 11*10^2+11 nevermind about that one

myininaya:

because it does have a quadratic form

myininaya:

that is factorable

myininaya:

i was thinking about other polynomials like x^4+x^3+x+1 when I wrote 11011

myininaya:

but yep it has form 11*x^2+11 which is 11(x^2+1)

myininaya:

anyways it is my bed time i have work tomorrow

myininaya:

thanks moreso for chatting with me i was bored

celticcat:



celticcat:

- thats a really old way to find square roots that works for all numbers

celticcat:

first you partition the number from the left into pairs then you find the nearest square root less than the first pair in this case its exactly 5 so you place 5 in the answer line (top line) and 5 to the left of vertical line 5*5 = 25 the subtract next double th e number in answer line and place it left of vertical line next bring down the next 2 digits (10) as shown now we want to add 1 digit to the 10 so that 10x <= 0010 in this case it is 0 so we add 0 to the answer line Now bring down the next 2 digits 00 so we have 001001 Now double the 50 on answer line and place it to the left to get 1001 we need to add 1 digit to the 100 so 1 is placed on answer line and we are finished we have the exact square root.

celticcat:

its long wined I know!! much easier with a calculator!

celticcat:

* long winded

myininaya:

That's interesting.

myininaya:

it kinda looks like division

celticcat:

I've tried to see if there's an algebraic basis for it - if we can reconcile it with the binomial (a + b^2 = a^2 + 2ab + b^2??

Hero:

What @celticcat presented is something more efficient, algorithmic, and convenient as an alternative to the calculator. Thanks myininaya for posting and thanks celticcat for your solution. I figured out how to do it based on your solution.

Hero:

For example, I figured out how to find the Square root of 66049 using Square Root Long Division Method: 1gVn4kPImo9pyD4wEQkmJmebG6ZMdv6W.png

totally17:

501

282006:

I got 501 What i did was divided it by two and then there is you answer

katwhirles:

501

darkknight:

501

smackz:

oh wow 3 years ago

Savannah23:

501

PrestonTX:

2345678765456789876543234567890987654321345678909876543234567890987654321234567890-098765432134567890987654345678909876543456789876543456789876543456789876456789876543223456789876543234567899876543234567899876543234567898765434567898765434567890987654323456789876543234567890987654321234567890987654323456789098765432

blubabub:

The square root of -251001 is the number, which multiplied by itself 2 times, is -251001. In other words, this number to the power of 2 equals -251001

5StarFab:

Ain't this is the wrong section

Hero:

@5starfab wrote:
Ain't this is the wrong section
What were you even doing to find this I wonder.